# https://www.outerlands.io/

### v2

@@ Abstract
Optimistic rollups are a popular and promising method of increasing the throughput capacity of their underlying chain. These methods rely on economic incentives to guarantee their security. We present a model of optimistic rollups that suggests that the incentives are not necessarily aligned with the expected behavior of the players, thus potentially undermining the security of existing optimistic rollups. We discuss some potential solutions illuminated by our model. ollup, Blockchain, Scaling
@@ 1 Introduction
Blockchains are public ledgers used to record transactions and execute decentralized programs. Rollups are a ‘Layer 2’ (L2) scaling solution that operate on top of a main blockchain [16]. They add throughput capacity by moving the computations necessary for transactions off the mainchain on to the L2 chain. The data is still posted on the mainchain; a fact that constitutes the main difference between rollups and other scaling solutions. There are two major types of rollups: zero knowledge rollups and optimistic rollups. Optimistic rollups function by ‘optimistically’ expecting that the commit transactions to a blockchain is being done accurately and honestly in general [16, 5]. The execution of these commitments is overseen by other nodes, but they are not necessarily systematically checked. These differ from zero knowledge rollups, which require every commitment to be proved correct every time.
@@
Optimistic rollups function by having two types of nodes. There are nodes that collect and perform the calculations for the bundle of transactions. They must put up a stake to participate. We will call these nodes aggregators, but note that this role is sometimes also called the sequencer. The work done by these nodes is then checked by other nodes that watch the chain and submit fraud proofs when they suspect an inaccuracy. This second role is called a validator. These validators must make a stake before they can press a challenge and they get rewarded if the challenge is successful and penalized if it is not. If a challenge is successful, the validator receives some of the aggregator’s stake and the rest is burned. If a fraud proof is brought, it must be settled on the mainchain, which can be costly, so the system is intended to disincentivize unnecessary proofs [16, 5].
@@
Currently, the most popular implementations of optimistic rollups are Arbitrum and Optimism, with approximately 65% and 19% of the market share, respectively [12]. They differ in their precise implementation of the fault proof procedure. Arbitrum has a multi-step procedure that is designed to lessen the load on the main chain, although it still retains the same staking mechanism. Optimism has a simple procedure for resolving fault proofs, but it is more onerous on the main chain and thus more expensive. Neither implementation has decentralized aggregator, although they both state in their roadmap that decentralization is a priority and coming soon. Moving forward, we will assume that the aggregator is a decentralized and not a trusted party, but will remain agnostic on how precisely the fault proof is executed.
@@
Optimistic rollups require users to wait to withdraw their deposits until a ‘challenge period’ has ended. This allows time for the validator nodes to check the transactions. This period is usually seven days. There are proposed solutions to this issue, for example the use of liquidity pools to allow users to access funds before their transactions are final, but the waiting period will always be an inherent part of the protocol.
@@ 1.1 Related Work
A concurrent work by Li[9] engages with a similar question. In the working paper, the author presents a model of a potential attack on the well functioning of optimistic rollups and concludes that their current design is not secure. Here, we consider a mechanism to be secure only if there is a pure equilibrium that is the intended behaviour. There are some substantial differences between their model and ours. We assume that the choice on the part of the validator of whether or not to perform the search is distinct from the choice of whether or not to challenge, i.e. a validator can blindly challenge without having performed a search. We also assume different payoffs. If a dishonest transaction is caught, we assume the aggregator does not get the utility of the attack, that is, the faulty commitment is revoked. So a caught dishonest aggregator has their stake slashed while also losing the utility of the attack. Furthermore, we assume that an erroneous challenge results in a penalty to the validator.
While Li’s model and the model we present are different, the general conclusion, that optimistic rollups are not incentive compatible, is similar. We believe that our model captures the phenomenon in a more nuanced and accurate way. The paper by Li [9] also extends their model to a case with multiple of each type of player. A similar analysis with many players is was done by researchers at Offchain Labs in another concurrent work. They also find an undesirable equilibrium when there is only one of each player [10].
Other works have used rational and economic augments to investigate other blockchain designs. This type of analysis has been applied to Byzantine Fault Tolerance [1, 7], proof-of-stake [14, 15], sharding [11], and frequently to proof-of-work systems [8, 3, 6]. A survey paper details many different applications of game theory to blockchain scaling mechanisms [16].
There has also been a great deal of research into the improvement and functioning of scaling solutions [4]. The work in Thibault et al [16] discusses many types of blockchain scaling solutions and we follow their account of the functioning of optimistic rollups in the setup of our model.
@@ 2 Our Model
Our model is set up with three players: the aggregator $A$, the validator $V$, and the transactor $T$. The transactors are the owners of the transactions that are to be committed. These are not necessarily transactions; they may instead be part of the use of a smart contract. The transactor reaps some utility $u_{T}$ from the execution of their transaction and pays fee $f$, which goes to the aggregator as compensation for their services. We assume that the utility of the transaction is greater than the fee, i.e. $u_{T}>f$. In order to commit the transaction, the aggregator must perform some underlying computational work that incurs cost, $w$. This is a computational cost, not a monetary cost, but we will assume that this and other computational costs are easily transformed into monetary costs. We will assume that the agents enjoy simple linear utility. Thus, in this stylized model of a transaction, we have the following utilities, if all goes well:
$A:$ $\quad f-w$
$V:$ 00$\quad$ 0
$T:$ $\quad u_{T}-f$
The algorithm functions as follows.
Algorithm 2.1
Protocol.
  • Transactor $T$ sends their payment $f$ and their transaction information to the aggregator $A$.
@@
        If the proof is found to be correct, the payment and stake can be withdrawn by $A$ and $V$’s stake is slashed.
        If the proof is found to be incorrect, $T$ is able to withdraw their $f$ and half the aggregator stake $s_{V}$ is given to $V$ and the other half is burned.
@@
Algorithm 1: Optimistic rollup
With this algorithm in hand, we can start to look at the incentives for each party. Initially, we will assume that the validator nodes incur no cost to keep an eye on the activities on chain, but we will introduce this search cost in the next section.
One important question is what happens when something goes wrong. If a lazy aggregator acts dishonestly, they might just abscond with the fee $f$. Rather than being content with just the fee, the aggregator could siphon currency from a smart contract. This need not come from the transactor, but the ecosystem at large. A particularly large sum might be siphoned from a bridge contract, which is how the L2 connects to the L1 or another chain. One attack, though different means than those detailed here, was able to steal \$326M from a Solana bridge contract [13]. We call this potentially large benefit $z$. Attempting this attack will require about the same computational work $w$. With these assumptions, we can readily write down the normal form game and find the equilibrium.
@@
The optimistic rollup paradigm as described in Algorithm 1 yields ($P$ behaving honestly) and ($V$ posing no challenge) as the pure-strategy Nash equilibrium when $V$ has perfect information and all players are rational and risk neutral.
@@
We have the game in extensive form as follows, with utility vectors of the form $\begin{pmatrix}A\\ V\\ T\end{pmatrix}$:
@@
Given these strategies for $V$, $A$ can see that honest behaviour will result in a utility of $f-w$, which is positive under our assumptions, and dishonest behaviour will result in a utility of $-s_{P}$. Thus, $A$’s best strategy is to behave honestly. The transactor will receive $u_{T}-f$, which is assumed positive, and will therefore not chose the outside option. Thus ($A$ behaving honestly) and ($V$ posing no challenge) is the pure-strategy Nash equilibrium, as desired. ∎
If our assumptions were true to reality, this finding would support the incentive compatibility of optimistic rollups. We will now discuss the roles of search costs.
@@ 3 Search costs
Even without introducing search costs, it is easy to read off the fact that, given that honest behavior on the part of the prover is rational, $V$ will have an expected utility of 00. Why then would a validator participate in the protocol? We observe participation in real implementations and there are good reasons to participate other than the potential of the reward. These include the fact that the validator nodes might be run by people with an interest in the rollup protocol. Frequent users, for example, might want to keep an eye on the blockchain in order to be assured of the security of their own transactions. Running a node gives access to the shared ledger, which grants access to data that is important for decentralized applications or frequent traders. A simpler solution purposed by Arbitrum is to just pay validator nodes for their services [2].
Running a validator node is not as computationally intensive as other types of nodes in other applications. In order to model this accurately, we introduce a small search cost. If the validator incurs this search cost, they have perfect information about $A$’s honesty.
In order to model this addition, we introduce another action where the validator decides whether they intend to invest in the search cost or not. For simplicity, we consider this to be a binary choice. We also omit the transactors $T$ in this game for simplicity because they are not in a position to make a choice in the game. They could opt for an outside option, but we will consider this once we have explored the game with just $A$ and $V$.
In the following, the dotted lines represent an information sets. That is, the player does not know which of the two nodes they are really at. In context, $A$ will not know if $V$ is performing searches or not, and, if $V$ has chosen (No Search), they do not know if $A$ is honest.
In order to simplify the calculations, we normalize the honest behavior of the aggregator to 00. That is set $f-w=0$. We will also simplify the penalty that $A$ faces when they misbehave to be just the relatively large slashed stake $s_{A}$ and disregard the $w$ term.
The utility vectors are of the form $\begin{pmatrix}A\\ V\end{pmatrix}$.
@@
The right tree corresponds to the game we have already discussed, except that $A$ does not know that $V$ is performing searches. We now assign three probabilities to the choices of the players: $V$ does no search with probability $b$ and then challenges with probability $g$; and $A$ is honest with probability $h$. All are between 00 and $1$.

### v3

# Incentive Non-Compatibility of Optimistic Rollups

@@ Abstract
Optimistic rollups are a popular and promising method of increasing the throughput capacity of their underlying chain. These methods rely on economic incentives to guarantee their security. We present a model of optimistic rollups that shows that the incentives are not aligned with the expected behavior of the players, thus potentially undermining the security of existing optimistic rollups. We discuss some potential solutions illuminated by our model. ollup, Blockchain, Scaling
@@ 1 Introduction
Blockchains are public ledgers used to record transactions and execute decentralized programs. Rollups are a ‘Layer 2’ (L2) scaling solution that operate on top of a main blockchain [scaling_survey]. They add throughput capacity by moving the computations necessary for transactions off the mainchain on to the L2 chain. The data is still posted on the mainchain; a fact that constitutes the main difference between rollups and other scaling solutions. There are two major types of rollups. Optimistic rollups function by ‘optimistically’ expecting that transactions are being committed to a blockchain accurately and honestly in general [scaling_survey, op]. The execution of these commitments is overseen by other nodes, but are not necessarily systematically checked. These differ from zero knowledge rollups, which require every commitment to be proved correct every time.
@@
Optimistic rollups function by having two types of nodes. There are nodes that collect and perform the calculations for the bundle of transactions. They must put up a stake to participate. We will call these nodes aggregators, but note that this role is sometimes also referred to as the sequencer. The work done by these nodes is then checked by other nodes, validators that watch the chain and submit fraud proofs when they suspect an inaccuracy. These validators must stake currency before they can press a challenge and lose their stake if this challenge is erroneous. If a challenge is successful, the validator receives some of the aggregator’s stake and the rest is burned. If a fraud proof is brought, it must be settled on the mainchain, which can be costly, so the system is intended to disincentivize unnecessary proofs [scaling_survey, op].
@@
Currently, the most popular implementations of optimistic rollups are Arbitrum and Optimism, with approximately 65% and 19% of the market share, respectively [mourya_ethereum_2023]. They differ in their precise implementation of the fraud proof procedure. Neither implementation has a decentralized aggregator, although they both state in their roadmap that decentralization is a priority and coming soon. Moving forward, we will assume that the aggregator is decentralized, but will remain agnostic on how precisely the fault proof is executed.
@@
Optimistic rollups require users to wait to withdraw their deposits until a ‘challenge period’ has ended. This allows time for the validator nodes to check the transactions. This period is usually seven days. There are proposed solutions to this issue, for example the use of liquidity pools to allow users to access funds before their transactions are final, but the waiting period will always be an inherent part of the protocol. A more pressing and well-known problem is that of the ‘verifier’s dilemma’, which observes that the verifier has no incentive to execute its role if there are never any errors and therefore no rewards. In this paper, we present a formal model of the verifier’s dilemma and discuss potential solutions that our model illuminates.
@@ 1.1 Related Work
The verifier’s dilemma is a frequently discussed issue, with much of that discussion happening in informal blogs. One of the aims of this paper is to present a formal model of an already well-established line of informal inquiry. A concurrent work by Li[li] engages with a similar question. In the paper, the author presents a model of a potential attack on the well functioning of optimistic rollups and concludes that their current design is not secure. Here, we consider a mechanism to be secure only if there is a unique pure equilibrium that is the intended behaviour. There are some substantial differences between their model and ours. We assume that the choice on the part of the validator of whether or not to perform the search is distinct from the choice of whether or not to challenge, i.e. a validator can blindly challenge without having performed a search. We also assume different payoffs. If a dishonest transaction is caught, we assume the aggregator does not get the utility of the attack, that is, the faulty commitment is revoked. So a caught dishonest aggregator has their stake slashed while also losing the utility of the attack. Furthermore, we assume that an erroneous challenge results in a penalty to the validator.
While Li’s model and the model we present are different, the general conclusion, that optimistic rollups are not incentive compatible, is similar. We believe that our model captures the phenomenon in a more nuanced and accurate way. The paper by Li [li] also extends their model to a case with multiple of each type of player. A similar analysis with many players was done by researchers at Offchain Labs in another concurrent work. They also find an undesirable equilibrium when there is only one of each player [mamageishvili]. Our model alone captures the role of search costs and the ability to randomly challenge. We show that the desired behaviour is not incentive compatible, even with only one of each player. This work shows the gravity of the verifier’s dilemma in a succinct but accurate model, which constitutes an important tool the community has lacked.
Other works use rational and economic augments to investigate other blockchain designs. This type of analysis has been applied to Byzantine Fault Tolerance [bft2, citation-key], proof-of-stake [saleh, eth_three_attacks], sharding [sharding], and proof-of-work systems [pow, btc, btc_2]. A survey paper details diverse applications of game theory to blockchain scaling mechanisms [scaling_survey]. There is extensive research into the improvement and functioning of scaling solutions [committee_scaling]. The work in Thibault et al [scaling_survey] discusses many types of blockchain scaling solutions and we follow their account of the functioning of optimistic rollups in the setup of our model.
@@ 2 Our Model
Our model is set up with three players: the aggregator $A$, the validator $V$. The owners of the transactions are not taken into account, but they pay some fee $f$, which goes to the aggregator as compensation. In order to commit the transaction, the aggregator must perform some underlying computational work that incurs cost, $w$. This is a computational cost, not a monetary cost, but we will assume that this and other computational costs are easily transformed into monetary costs. We will assume that the agents enjoy simple linear utility. Thus, in this stylized model of a transaction, we have the following utilities, if all goes well: $A:f-w$ and $V:0$. The algorithm functions as follows.
Algorithm 1
Optimistic Rollup Protocol.
  1. 1.
    Transactor sends their payment $f$ and their transaction information to the aggregator $A$.
  2. 2.
@@
      1. 4.1.1.
        If the challenge is found to be erroneous, the payment and stake can be withdrawn by $A$ and $V$’s stake is slashed.
      2. 4.1.2.
        If the challenge is found to be correct, half the aggregator stake $s_{A}$ is given to $V$ and the other half is burned.
    2. (4.2)
@@
With this algorithm in hand, we can begin clarifying the incentives for each party. Initially, we will assume that the validator nodes incur no cost to keep abreast of the activities on chain, but we will introduce this cost in the next section. Thus, we are assuming that $V$ always searches for free, and therefore has perfect information.
One important question is what happens when something goes wrong. If a lazy aggregator acts dishonestly, they might just abscond with the fee $f$. Rather than being content with just the fee, the aggregator could siphon currency from a smart contract. A particularly large sum might be siphoned from a bridge contract. One attack, through different means than those detailed here, was able to steal \$326M from a Solana bridge contract [bridge]. We call this potentially large benefit $z$. Attempting this attack will require about the same computational work $w$. With these assumptions, we can readily write down the extensive form game and find the equilibrium.
@@
The optimistic rollup paradigm as described in Algorithm 1 yields ($A$ behaving honestly) and ($V$ posing no challenge) as the pure-strategy Nash equilibrium when $V$ has perfect information and all players are rational and risk neutral.
@@
We have the game in extensive form as follows, with utility vectors of the form $(A,V)^{T}$:
@@
Given these strategies for $V$, $A$ can see that honest behaviour will result in a utility of $f-w$, which is positive under our assumptions, and dishonest behaviour will result in a utility of $-s_{P}$. Thus, $A$’s best strategy is to behave honestly. Thus ($A$ behaving honestly) and ($V$ posing no challenge) is the pure-strategy Nash equilibrium, as desired. ∎
@@ 3 Search costs
Even without introducing search costs, it is easy to read off the fact that, given that honest behavior on the part of the prover is rational, $V$ will have an expected utility of 00. Why then would a validator participate in the protocol? We observe participation in real implementations and there are good reasons to participate other than the potential of the reward. These include the fact that the validator nodes might be run by parties with an interest in the rollup protocol. Frequent users, for example, might want to keep an eye on the blockchain in order to be assured of the security of their own transactions. Running a node gives access to the shared ledger, which grants access to data that is important for decentralized applications or frequent traders. Running a validator node is not as computationally intensive as other types of nodes in other applications. In order to model this accurately, we introduce a small search cost, $x$. If the validator incurs this search cost, they have perfect information about $A$’s honesty.
In order to model this addition, we introduce another action where the validator decides whether they intend to invest in the search cost or not. In the following, the dotted lines represent information sets. In context, $A$ will not know if $V$ is performing searches or not, and, if $V$ has chosen (No Search), they do not know if $A$ is honest. In order to simplify the calculations, we normalize the honest behavior of the aggregator to 00. That is set $f-w=0$. We will also simplify the loss that $A$ faces when they misbehave to be just the relatively large slashed stake $s_{A}$ and not the $w$ term.
The utility vectors are of the form $(A,V)^{T}$.
@@
The right tree corresponds to the game we have already discussed, except that $A$ does not know that $V$ is performing searches. We now assign three probabilities to the choices of the players: $V$ does no search, and makes choices blindly, with probability $b$ and then challenges with probability $g$; and $A$ is honest with probability $h$. All are between 00 and $1$. Thus we can read off the utility functions, starting with $A$. If the aggregator is honest, they will receive 00 utility, regardless of $V$’s actions. Recall this is due to our normalization of $f-w$ to 00. When $V$ is blind and is challenging with a probability $g$, dishonesty results in utility $g(-s_{A})+(1-g)z$. When $V$ is not blind, they challenge if and only if $A$ is indeed dishonest, yielding $-s_{A}$ utility for $A$. Combining these pieces, we get,
$A:\quad h\cdot 0+(1-h)(b\cdot(g\cdot(-s_{A})+(1-g)z)+(1-b)(-s_{A}))$ (1)
When $V$ chooses no search, we can see that the validator reaps 00 utility if they do not challenge. In the challenge case, if $A$ is honest, $V$ has their stake burned for utility $-s_{V}$. In the dishonest case, they receive $\frac{1}{2}s_{A}$. Weighting this with the probabilities, we find the left hand side has expected utility $g\cdot\left(h\cdot(-s_{V})+(1-h)\frac{1}{2}s_{A}\right)$. On the right hand side, $V$ can now make an informed decision. If $A$ is honest, $V$ will not challenge, resulting in utility $-x$ and if $A$ is dishonest, $V$ will always challenge for utility $\frac{1}{2}-x$. Weighting these terms by the honesty probability yields the right hand side utility of $h\cdot(-x)+(1-h)\left(\frac{1}{2}s_{A}-x\right)$, which can be simplified to $(1-h)\left(\frac{1}{2}s_{A}\right)-x$. To get the whole expected utility, we weight both sides of the tree by the probability of no search $b$, which gives,
$V:\quad bg\cdot\left(h\cdot(-s_{V})+(1-h)\frac{1}{2}s_{A}\right)+(1-b)\left((1-h)\frac{1}{2}s_{A}-x\right)$ (2)
3.1 Main Result
It suffices to show that there exist values for which the equilibrium is not the desired outcome of full search and honesty. The fully general proof is available in the online version, but the example here more readily illuminates the key interactions. We start by noting that we should have the following ordinal constraint based on our context: $x<s_{V}\leq s_{A}\ll z$. Take $s_{A}=s_{V}=s$, $x=\frac{s}{24}$, and $z=24s$, convenient choices among many that clearly obey these constraints.
Lemma 2
A rational, risk neutral $A$ will be indifferent to being honest if $s_{A}=s_{V}=s$ and $z=24s$ and $g=1-\frac{1}{25b}$.
Proof
From Eq. 1, we can see that dishonest behavior results in utility of $b\cdot(g\cdot(-s)+(1-g)24s)+(1-b)(-s)$, which is easily simplified to $25bs-25bgs-s$. Since we want to find the values of $b$ and $g$ for which $A$ is indifferent between honesty and dishonesty, we set this expression equal to 0, the utility of honesty. We solve for $g$ first, readily getting,
$g=1-\frac{1}{25b}$ (3)
Given that $g$ must be positive, we can read off that $b>\frac{1}{25}$. ∎
Theorem 3.1
Given $s_{A}=s_{V}=s$, $x=\frac{s}{24}$, and $z=24s$, there are values of $b$, $g$, and $h$ that admit a mixed equilibrium that is not the same as that found in Lemma 2, if both $A$ and $V$ are rational and risk neutral.
Proof
From Eq. 3, we can see that a value of $b=\frac{1}{5}$ yields $g=\frac{4}{5}$ and both are viable. Given these values, we can ask, for which value $h$, determined by $A$, is $V$ indifferent. We can now simply plug these values into Eq. 2 and set it to 0, yielding,
$\frac{1}{5}\left(\frac{4}{5}\left(h\left(-s\right)+\left(1-h\right)\frac{s}{2}\right)\right)+\frac{4}{5}\left(\left(1-h\right)\frac{s}{2}-\frac{s}{24}\right)=0$
Solving for $h$ yields $h=\frac{67}{96}$. Thus, we have found values for which both $A$ and $V$ can force the other to be indifferent to their choice of behavior. Thus we have found an equilibrium not in keeping with the desired functioning. ∎
4 Potential Solutions
The verifier’s dilemma has yet to become a practical problem because existing optimistic rollups have yet to decentralize their sequencers. While this is one potential solution, it is important to ask how decentralization might be achieved without compromising the security of the chain. Another trivial solution would be to revert to a mechanism that requires the validators to check the integrity of the block every time, as is done with proof of stake L1s and zero knowledge rollups. We present here an outline of potential solutions that are more in keeping with the spirit of optimistic rollups.
4.1 Random Checks
If we were to stipulate that each block be checked, we would stray away from the spirit of optimistic rollups and lose their lightweight efficiency. Requiring intermittent, random checks could keep the optimistic efficiency without the hazard of the verifier’s dilemma. Whether such a set up would indeed be meaningfully more efficient than simply checking every time would depend on the relative value of the potential benefit from attack and the potential penalties. It would also depend on precisely how the random checks would be implemented. This would not necessitate fraud proofs, which are needed to find precisely where the chain became faulty, and instead would require proofs of just the block in question. These proofs could be done by nodes, which would again need incentives, or by smart contracts on the L1, which could have expensive execution that might fall on the users or be paid for with inflation. If we assume this could be done with a smart contract without too much additional cost to the user, we get a game where $C$ is the contract ‘player’ in the place of $V$. Let us also define $p$, where $10<p<$ 1, to be the probability with which the contract checks the proof.
Lemma 3
There are values of the probability of check, $p$, $s_{A}$, and $z$ such that a rational, risk neutral aggregator will always be honest.
{istgame}
[…]
Game 3: Game tree illustrating a random check and the utility of $A$.
It is immediate that the expected income on the left child of the root is zero. Recall this is because we normalized it so. Since the left, honest choice on the part of $A$ is the desired outcome, we need a $p$ such that the right side yields negative income in expectation. Thus we would need, $(1-p)z-ps_{A}<0$. Solving for $p$ readily yields $p>\frac{z}{z+s_{A}}$. This is clearly in keeping with our requirement that $10<p<$ 1, but is also clear that, given $s_{A}\ll z$, $p$ will be close to $1$. Thus, there are only minor savings as compared to checking every time, but there are values for which such a setup would result in a rational, risk neutral $A$ behaving honestly, as desired. ∎
4.2 Random Mistakes
4.3 Easter Eggs
If the search costs, $x$, are relatively small, compensating the validators directly for their services seems like a tempting solution, and has been considered by Arbiturm [arbitrum]. However, it may not be clear whether or not the search is being performed correctly and checking this search would be yet another role with its own incentives. If it were possible to hide reward ‘Easter eggs,’ which could be found through proper searching but not a random challenge, this could act as a delivery mechanism for validator rewards. These rewards would be realized even when the aggregator behaves honestly.
Lemma 4
Compensating the validator with some term with expected value $y\geq x$ will result in the expected income for the validator of always searching being higher than not searching.
Let us add $y$, the expected value of finding the Easter egg, to the validator income in the case of search, i.e. when $b=0$, which we can read off from Eq. 2. This should yield greater expected utility than never searching, i.e. when $b=1$.
$(1-h)\frac{1}{2}s_{A}-x+y>-ghs_{V}+g(1-h)\frac{1}{2}$
since $g<1$, it is clear to see that if we have $-x+y>-ghs_{V}$, the above term is always true. If we require that $y\geq x$, the result is immediate. ∎
[…]
Appendix